Find two consecutive odd integers such that six times the first equals ten more than four times the second.
Answer by checkley77(12844)

Let x & (x+2) be the 2 consecutive odd integers.
6x=4(x+2)+10
6x=4x+8+10
6x-4x=18
2x=18
x=18/2
x=9
ans for the smaller integer.
9+2=11 ans. for the next integer.
Proof:
6*9=4*11+10
54=44+10
54=54